Substitution: 2x-3y=1 and 3x+2y=8
Do help yay

Answer by mananth(16946)

2 x + -3 y = 1 .............1
-3 y = 1 + -2 x
/ -3 =
y = 0 + 1 x
3 x + 2 y = 8 .....................2
Substitute y in (2)
3 x + 2 ( 0 + 1 x
3 x + -1 + 1 x = 8
3 x 1 1/3 x = 8 2/3
4 1/3 x = 8 2/3
/ 4 1/3
x= 2
Plug the value of x in (1)
2 x + -3 y = 1
2 * 2 + -3 y = 1
4 + -3 y = 1
-3 y = -3
/ -3
y= 1
x=2,y=1
m.ananth@hotmail.ca